A bicyclist rode into the country for 5 hours. In returning, her speed was 5 miles per hour faster and the trip took 4 hours. What was her speed each way?
Answer by checkley71(8403)

5X=4(X+5) OR 5X=4X+20 OR X=20 THE SPEED GOING & 20+=25 PROOF
5*20=4(20+5) OR 100=4*25 OR 100=100